OpenStudy (anonymous):

How would I find the derivative sin^-1(x^2)? Please I've been at this for a long time

OpenStudy (anonymous):

It's the derivative of \[\sin ^{-1}(x^2)\] I know \[\sin y= x ^{2}\]

OpenStudy (solomonzelman):

yes, then differentiate on both sides

OpenStudy (solomonzelman):

y' cos(y) = 2x

OpenStudy (solomonzelman):

then cos(y)= sqrt( 1- sin^2(y) )

OpenStudy (solomonzelman):

y ( sqrt (1 - x^4 ) ) = 2x

OpenStudy (solomonzelman):

I meant y' ( sqrt (1 - x^4 ) ) = 2x

OpenStudy (solomonzelman):

y' = 2x/ ( sqrt (1 - x^4 ) )

OpenStudy (anonymous):

oaky thank you!